# Tensor Completion via Integer Optimization

### v1

@@ Abstract
The main challenge with the tensor completion problem is a fundamental tension between computation power and the information-theoretic sample complexity rate. Past approaches either achieve the information-theoretic rate but lack practical algorithms to compute the corresponding solution, or have polynomial-time algorithms that require an exponentially-larger number of samples for low estimation error. This paper develops a novel tensor completion algorithm that resolves this tension by achieving both provable convergence (in numerical tolerance) in a linear number of oracle steps and the information-theoretic rate. Our approach formulates tensor completion as a convex optimization problem constrained using a gauge-based tensor norm, which is defined in a way that allows the use of integer linear optimization to solve linear separation problems over the unit-ball in this new norm. Adaptations based on this insight are incorporated into a Frank-Wolfe variant to build our algorithm. We show our algorithm scales-well using numerical experiments on tensors with up to ten million entries.
Tensor Completion, Optimization
@@ 1 Introduction
A tensor is a multilinear operator, and it can be represented as an array of numbers referenced by multiple indices. Though vectors and matrices are special cases of tensors, tensors with three or more indices pose unique challenges. Many tensor problems are np-hard (Hillar & Lim, 2013), such as computing: rank, singular values, and nuclear norm.
In this paper, we address one such difficult problem called tensor completion. Here, a small subset of tensor entries are observed – possibly with noise. Under an assumption of low-rankness, the problem is to fill-in the remaining, unobserved entries – and remove noise, if any. Since modern datasets are often multidimensional, there are many applications of tensor completion (Song et al., 2019), including: recommendation systems (Ge et al., 2016; Karatzoglou et al., 2010), information diffusion (Zafarani et al., 2014), regression (Aswani, 2016), computer vision (Duarte & Baraniuk, 2012; Signoretto et al., 2011; Tan et al., 2013), and bioinformatics (Bazerque et al., 2013; Acar et al., 2011).
@@ 1.1 Related Work
The information-theoretic rate for estimation error is $\sqrt{k\cdot\sum_{i}r_{i}/n}$ for a tensor completion problem, where: $k$ is tensor rank, $r_{i}$ is the $i$-th dimension of the tensor, and $n$ is the number of samples (Gandy et al., 2011). Previous works have tried to characterize the tradeoff between complexity and the information-theoretic rate. For example, Barak & Moitra (2016) uses the Rademacher complexity of a sum-of-squares hierarchy to suggest a suitable norm for tensor completion, and observe a gap between what can be achieved information theoretically and what is attained by their computationally efficient method.
Initial work on tensor completion used decomposition methods based on CP and Tucker decomposition (Tomasi & Bro, 2005). Other approaches accounted for robustness to outliers and data corruptions (Javed et al., 2015; Jain et al., 2017), or imposed fixed-rank constraints (Kressner et al., 2014). Alternative approaches used various tensor norms as a convex surrogate for tensor rank (Montanari & Sun, 2018; Yuan & Zhang, 2016; Barak & Moitra, 2016). Our approach falls into this category, but here we define a novel norm.
Past approaches either traded information-theoretic rate for a computationally efficient algorithm (Montanari & Sun, 2018; Barak & Moitra, 2016); likewise, heuristics have been developed to compute non-certifiably-optimal solutions to such np-hard problems whose optimal solutions (if found) in principal achieve the information-theoretic rate (Kressner et al., 2014; Yuan & Zhang, 2016; Jain et al., 2017). In contrast, we present a globally convergent algorithm that attains the information-theoretic rate (hence data efficiency), while certifying optimal solutions on instances on general tensors of sizes up to $10^{\times 7}$ within minutes.
Certain algorithms for special cases of tensor completion have achieved the information-theoretic rate through practical computation. Aswani (2016) studied tensor completion for rank-1 non-negative tensors, formulating it as conic optimization with exponential constraints. Symmetric tensor completion was also studied, where Rao et al. (2015) used a variant of the Frank-Wolfe algorithm and Cai et al. (2019) used two-stage non-convex optimization. In our previous work (Bugg et al., 2022), we proposed an algorithm based on first-order optimization, that uses integer optimization for a weak separation oracle, for the special case of nonnegative tensors. We also explored accelerated versions of this integer-optimization-based algorithm for nonnegative tensors in (Pan et al., 2023).
@@ 1.2 Contribution
In this paper, we design an algorithm for tensor completion of general tensors, which converges to a global optimum with a linear number of oracle calls and satisfies the information-theoretic rate. In past literature, tensor completion algorithms that perform both information-theoretically and computationally-well have been achieved only in special cases (e.g., nonnegative tensors, symmetric orthogonal tensors).
Our algorithm is a step in this direction for the case of general tensors. Building upon applications of various special-case tensor completion algorithms (Li et al., 2008; Aswani, 2016; Liu et al., 2012; Zhang et al., 2019), our algorithm for tensor completion opens up new applications that require positive and negative entries. Examples include social computing (Song et al., 2019), the moment method on multivariate distribution (Montanari & Sun, 2018), and healthcare applications (Gandy et al., 2011; Bazerque et al., 2013).
@@
Consequently, the tensor completion problem using gauge norm is also np-hard to solve to arbitrary accuracy. Nevertheless, since the formulation is a convex optimization, we design an algorithm using Blended Conditional Gradients (BCG) (Braun et al., 2019). We construct a weak separation oracle for BCG using an integer linear optimization formulation and an additional heuristic to accelerate computation. Our numerical experiments demonstrate that this algorithm achieves the information-theoretic rate and is efficient for tensors with as large as ten million entries.
@@ 1.3 Going from Nonnegative to General Tensors
Since the approach in our previous work (Bugg et al., 2022) is closely related to this paper, here we provide a brief discussion about why going from nonnegative to general tensors requires the design and analysis of a new tensor completion algorithm:
The key idea in our previous work was to define a gauge norm using a 00-$1$ polytope that represents the convex hull of all rank-1 nonnegative tensors whose maximum entry is 1. The polytope was designed in such a way that linear separation problems over the polytope could be written using integer linear constraints. The approach can be illustrated by an example: For the set $\{\zeta=\prod_{k=1}^{p}\nu_{k}:\nu_{k}\in\{0,1\}\}$ defined using a product, its linearization is given by $\{\zeta:0\leq\zeta\leq\nu_{k},\sum_{k}\nu_{k}+(1-p)\leq\zeta,\nu_{k}\in\{0,1\}\}$. A direct linearization is possible because of the properties that: $\zeta=0$ if any single $\nu_{k}=0$, and $\zeta=1$ if all $\nu_{k}=1$.
The natural generalization of this idea, which we pursue in this paper, is to design a gauge norm using a polytope whose vertices have ±$\pm 1$ entries and that represents the convex hull of all (general) rank-1 tensors whose maximum/minimum entry is ±$\pm 1$. However, the above property used for linearization no longer holds when considering the new polytope. This can be shown via an example: For the set $\{\zeta=\prod_{k=1}^{p}\nu_{k}:\nu_{k}\in\{-1,+1\}\}$, a direct linearization is not possible because $\zeta=+1$ if an even number of $\nu_{k}=-1$, and $\zeta=-1$ if an odd number of $\nu_{k}=-1$. The generalization is fundamentally different because it involves the parity (i.e., odd or even) of the underlying items being multiplied. This is much more challenging and requires new computational design and theoretical analysis.
General tensors also have well-posedness issues that do not occur for nonnegative tensors. In particular, the best low-rank approximation problem is not well-posed for tensors (de Silva & Lim, 2008); this is related to a well-known example that a sequence of rank-2 tensors can be constructed whose limit has rank-3. In contrast, the best low-rank approximation problem is well-posed for nonnegative tensors (Qi et al., 2016). These phenomena create a technical challenge in determining whether a new gauge norm as defined above can act as a convex surrogate for the rank of tensors.
@@ 2 Preliminaries
For a tensor $\psi\in\mathbb{R}^{r_{1}\times\cdots\times r_{p}}$ of order $p$, we use the notation $\psi_{x}:=\psi_{x_{1},\ldots,x_{p}}$ to refer to an entry corresponding to the indices $x=(x_{1},\ldots,x_{p})$, where $x_{i}\in[r_{i}]$ and $[s]:=\{1,\ldots,s\}$. Furthermore, we define some parameters that depend upon the tensor dimensions and order: $\rho=\sum_{i}r_{i}$, $\pi=\prod_{i}r_{i}$, and $\mathcal{R}=[r_{1}]\times\cdots\times[r_{p}]$.
By definition, a rank-1 tensor can be written as the tensor product of vectors, that is $\psi=\bigotimes_{k=1}^{p}\theta^{k}$, where $\theta^{(k)}\in\mathbb{R}^{r_{k}}$. Equivalently, each tensor entry is $\psi_{x}=\prod_{k=1}^{p}\theta^{(k)}_{x_{k}}$, where $\theta^{(k)}_{x_{k}}$ is the $x_{k}$-th element of vector $\theta^{(k)}$ for any index value $x_{k}\in[r_{k}]$. When obvious, we will use $\theta_{x_{k}}$ instead of $\theta^{(k)}_{x_{k}}$.
@@
$\operatorname{rank}(\psi)=\min\{q\ |\ \psi=\sum_{k=1}^{q}\psi^{k},\psi^{k}\in\mathcal{B}_{\infty}\text{ for }k\in[q]\}.$
Using this definition of tensor rank, a cp decomposition of the tensor is given by $\psi=\textstyle\sum_{k=1}^{\operatorname{rank}(\psi)}\psi^{k}$.
@@
where the constraint uses a new norm $\|\psi\|_{\pm}$ that we design below. We show that our norm does in fact act as a convex surrogate for tensor rank, analyze its complexity, and demonstrate that it leads to an efficient computation algorithm for the tensor completion problem.
@@ 3 Gauge Norm for Tensors
We construct a norm for general tensors using a gauge function (Chandrasekaran et al., 2012; Jaggi, 2013; Bugg et al., 2022). To start with, consider rank-1 tensors. Let $\mathcal{S}_{\lambda}$ be the set of rank-1 tensors such that each entry of the tensor has an absolute value of some $\lambda\in\mathbb{R}_{+}$:
@@
Our first step in defining a norm is to relate the convex hull of these sets of rank-1 tensors.
Proposition 3.1.
The convex hulls of these sets are the same, meaning we have $\mathcal{C}_{\lambda}:=\operatorname{conv}(\mathcal{B}_{\lambda})=\operatorname{conv}(\mathcal{S}_{\lambda})$.
Proof.
The proof for this proposition is similar to our proof of Proposition 2.1 in (Bugg et al., 2022), but where a multilinear optimization problem is formulated by restricting the entries to $[-1,1]$ instead of $[0,1]$. ∎
Remark 3.2.
@@ 3.1 Defining the Norm
Proof.
We use the result from Example 3.50 of (Rockafellar & Wets, 2009) to conclude that $\|\cdot\|_{\pm}$ is a norm. To apply the result, we check that the required conditions on $\mathcal{C}_{1}$ hold.
@@
The final, non-trivial condition required is that $\mathcal{C}_{1}$ has a non-empty interior. To prove this, we use Theorem 2.4 of (Rockafellar, 2015) that the dimension of $\mathcal{C}_{1}$ is the maximum of dimensions of simplices included in it. We construct a simplex in $\mathcal{C}_{1}$ that has dimension $\pi=\Pi_{i}r_{i}$, and hence $\mathcal{C}_{1}$ has a dimension of at least $\pi$. But the flattened vector space of tensors also has dimension $\pi$; consequently, the dimension of $\mathcal{C}_{1}$ is at most $\pi$. Hence, $\mathcal{C}_{1}$ has a dimension $\pi$, which is the full dimension of the space, implying that the set $\mathcal{C}_{1}$ must have a non-empty interior.
The rest of the proof constructs such a simplex of dimension $\pi$. Consider a polytope $D=\operatorname{conv}(0\>\cup\>\{d^{x}\}_{x\in\mathcal{R}})$. Here, for any $x=(x_{1},\cdots,x_{p})\in\mathcal{R}$, consider $d^{x}=\bigotimes_{k=1}^{p}\beta^{x_{k}}$ with each vector $\beta^{x_{k}}\in\mathbb{R}^{r_{k}}$ defined as
@@
With this, $\mathcal{C}_{1}$ is shown to satisfy all the required conditions, which means that the proposition holds. ∎
Towards our end goal of developing a practical algorithm for tensor completion using this gauge norm, our next result provides an alternative characterization of the vertices of $\mathcal{C}_{\lambda}$. This result is important because it shows that these vertices can be represented by linear inequality constraints using (binary) integer variables.
Proposition 3.4.
@@
Proof.
We show that the constraints defining the sets $\mathcal{S}_{\lambda}$ and $\widehat{\mathcal{S}}_{\lambda}$ are equivalent. Consider some $x\in\mathcal{R}$. From the definition of $\mathcal{S}_{\lambda}$ in (3), we have $\psi_{x}=\lambda\prod_{i=1}^{p}\theta_{x_{k}}$. Define $y_{x,k}=\prod_{i=k}^{p}\theta_{x_{k}}$, for $k\in[p]$ so that $\psi_{x}=\lambda y_{x,1}$. Or equivalently, in a recursive relationship, $y_{x,k}=\theta_{x_{k}}y_{x,k+1}$ for $k\in[p-1]$ and $y_{x,p}=\theta_{x_{p}}$. The recursive constraints can be thought of as a negated-XOR relation, and linearized by transformations for conjunctive and disjunctive statements (see Section 2.5 of (Conforti et al., 2014)). These linearized constraints correspond to constraints 2-5 in the definition of $\widehat{\mathcal{S}}_{\lambda}$ as given above. This shows that for each $x\in\mathcal{R}$, the tensor $\psi_{x}$ is defined the same in both $\mathcal{S}_{\lambda}$ and $\widehat{\mathcal{S}}_{\lambda}$. ∎

### v2

@@ Abstract
The main challenge with the tensor completion problem is a fundamental tension between computational effort and the information-theoretic sample complexity rate. Past approaches either achieve the information-theoretic rate but lack practical algorithms to compute the corresponding solution, or have polynomial-time algorithms that require an exponentially-larger number of samples for low estimation error. This paper develops a novel tensor completion algorithm that resolves this tension by achieving both provable convergence (in numerical tolerance) in a linear number of oracle steps and the information-theoretic rate. Our approach formulates tensor completion as a convex optimization problem constrained using a gauge-based tensor norm, which is defined in a way that allows the use of integer linear optimization to solve linear separation problems over the unit-ball in this new norm. Adaptations based on this insight are incorporated into a Frank-Wolfe variant to build our algorithm. We show our algorithm scales-well using numerical experiments on tensors with up to ten million entries.
Keywords:
Tensor Completion Gauge Norm Convex Optimization Franke-Wolfe
MSC:
15-06 90C10
@@ 1 Introduction
A tensor is a multilinear operator, and it can be represented as an array of numbers referenced by multiple indices. For a tensor $\psi\in\mathbb{R}^{r_{1}\times\cdots\times r_{p}}$ of order $p$, we use the notation $\psi_{x}:=\psi_{x_{1},\ldots,x_{p}}$ to refer to an entry corresponding to the indices $x=(x_{1},\ldots,x_{p})$, where $x_{i}\in[r_{i}]$ and $[s]:=\{1,\ldots,s\}$. Furthermore, we define some parameters that depend upon the tensor dimensions and order: $\rho=\sum_{i}r_{i}$, $\pi=\prod_{i}r_{i}$, and $\mathcal{R}=[r_{1}]\times\cdots\times[r_{p}]$. Though vectors and matrices are special cases of tensors, tensors with three or more indices pose unique challenges. Many tensor problems are np-hard hillar2013 , such as computing: rank, singular values, and nuclear norm.
In this paper, we address one such difficult problem called tensor completion. Here, a small subset of tensor entries are observed – possibly with noise. Under an assumption of low-rankness, the problem is to fill-in the remaining, unobserved entries – and remove noise, if any. Since modern datasets are often multidimensional, there are many applications of tensor completion song2019tensor , including: recommendation systems Ge_Caverlee_Lu_2016 ; Karatzoglou_Amatriain_Baltrunas_Oliver_2010 , information diffusion Zafarani_Abbasi_Liu_2014 , regression aswani2016 , computer vision Duarte_Baraniuk_2012 ; Plas_De_Moor_Suykens_2011 ; Tan_Cheng_Feng_Feng_Wang_Zhang_2013 , and bioinformatics Bazerque_Mateos_Giannakis_2013 ; Acar_Dunlavy_Kolda_Mørup_2011 .
@@ 1.1 Related Work
The information-theoretic rate for estimation error is $\sqrt{k\cdot\sum_{i}r_{i}/n}$ for a tensor completion problem, where: $k$ is tensor rank, $r_{i}$ is the $i$-th dimension of the tensor, and $n$ is the number of samples gandy2011tensor . Previous works have tried to characterize the tradeoff between complexity and the information-theoretic rate. For example, the approach in barak2016noisy uses the Rademacher complexity of a sum-of-squares hierarchy to suggest a suitable norm for tensor completion, and observe a gap between what can be achieved information theoretically and what is attained by their computationally efficient method.
Initial work on tensor completion used decomposition methods based on CP and Tucker decomposition Tomasi_Bro_2005 . Other approaches accounted for robustness to outliers and data corruptions Javed_Bouwmans_Jung_2015 ; Jain_Gutierrez_Haupt_2017 , or imposed fixed-rank constraints kressner2014low . Alternative approaches used various tensor norms as a convex surrogate for tensor rank montanari2018spectral ; yuan2016tensor ; barak2016noisy . Our approach falls into this last category, but we define a tensor norm using a gauge function as a convex surrogate.
Past approaches either traded information-theoretic rate for a computationally efficient algorithm montanari2018spectral ; barak2016noisy ; likewise, heuristics have been developed to compute non-certifiably-optimal solutions to such np-hard problems whose optimal solutions (if found) in principal achieve the information-theoretic rate kressner2014low ; yuan2016tensor ; Jain_Gutierrez_Haupt_2017 . In contrast, our algorithm is globally convergent and attains the information-theoretic rate (hence data efficiency), while certifying optimal solutions on instances on general tensors of sizes up to $10^{\times 7}$ within minutes.
Certain algorithms for special cases of tensor completion have achieved the information-theoretic rate through practical computation. Aswani aswani2016 studied tensor completion for rank-1 non-negative tensors, formulating it as conic optimization with exponential constraints. Symmetric tensor completion was also studied, where Rao et al rao2015 used a variant of the Frank-Wolfe algorithm and Cai et al cai2019nonconvex used two-stage non-convex optimization. Another algorithm based on first-order optimization, and that uses integer optimization for a weak separation oracle was proposed in Bugg et al bugg2022nonnegative for nonnegative tensors. Accelerated versions of this integer-optimization-based algorithm for nonnegative tensors were explored in Pan et al pan2023 .
@@ 1.2 Contribution
We design an algorithm for tensor completion of general tensors, which converges to a global optimum with a linear number of oracle calls and satisfies the information-theoretic rate. In past literature, tensor completion algorithms that perform both information-theoretically and computationally-well have been achieved only in special cases (some examples include nonnegative tensors, symmetric and orthogonal tensors).
Our algorithm is a step in this direction for the case of general tensors. Building upon applications of various special-case tensor completion algorithms li2008image ; aswani2016 ; liu2012tensor ; zhang2019robust , our algorithm for tensor completion opens up new applications that require positive and negative entries. Examples of applications of tensor completion include social computing song2019tensor , the moment method on multivariate distribution montanari2018spectral , and healthcare applications gandy2011tensor ; Bazerque_Mateos_Giannakis_2013 .
@@
Consequently, the tensor completion problem using gauge norm is also np-hard to solve to arbitrary accuracy. Nevertheless, since the formulation is a convex optimization, we design an algorithm using Blended Conditional Gradients (BCG) braun2019blended . We construct a weak separation oracle for BCG using an integer linear optimization formulation, for which only a linear number of calls are needed. Integer optimization has a collection of mature, commercial-grade solvers that can navigate complex trade-offs between compute and solution quality on NP-hard problems. Moreover, weak separation enables us to deploy an additional heuristic to practically accelerate the computation of oracle calls, as well as to apply early termination in the integer programming solver when a sufficiently good solution has been found. Our numerical experiments demonstrate that this algorithm achieves the information-theoretic rate and is efficient for tensors with as large as ten million entries.
@@ 1.3 Going from Nonnegative to General Tensors
Since the approach in Bugg et al bugg2022nonnegative is closely related to this paper, here we provide a brief discussion about why going from nonnegative to general tensors requires the design and analysis of a new tensor completion algorithm:
The key idea behind Bugg et al bugg2022nonnegative was to define a gauge norm using a 00-$1$ polytope that represents the convex hull of all rank-1 nonnegative tensors whose maximum entry is 1. The polytope was designed in such a way that linear separation problems over the polytope could be written using integer linear constraints. The approach can be illustrated by an example: Consider the set $\{\zeta=\prod_{k=1}^{p}\nu_{k}:\nu_{k}\in\{0,1\}\}$ defined using a product, its linearization is given by $\{\zeta:0\leq\zeta\leq\nu_{k},\sum_{k}\nu_{k}+(1-p)\leq\zeta,\nu_{k}\in\{0,1\}\}$. A direct linearization is possible in this case because of the properties that: $\zeta=0$ if any single $\nu_{k}=0$, and $\zeta=1$ if all $\nu_{k}=1$.
The natural generalization of the idea behind Bugg et al bugg2022nonnegative , which we pursue in this paper, is to design a gauge norm using a polytope whose vertices have ±$\pm 1$ entries and that represents the convex hull of all (general) rank-1 tensors whose maximum/minimum entry is ±$\pm 1$. However, the above property used for linearization no longer holds when considering the new polytope. This can be shown via an example: For the set $\{\zeta=\prod_{k=1}^{p}\nu_{k}:\nu_{k}\in\{-1,+1\}\}$, a direct linearization is not possible because $\zeta=+1$ if an even number of $\nu_{k}=-1$, and $\zeta=-1$ if an odd number of $\nu_{k}=-1$. The generalization is fundamentally different because it involves the parity (i.e., odd or even) of the underlying items being multiplied. This is much more challenging and requires new computational design and theoretical analysis.
General tensors also have well-posedness issues that do not occur for nonnegative tensors. In particular, the best low-rank approximation problem is not well-posed for tensors desilva2008 ; this is related to a well-known example that a sequence of rank-2 tensors can be constructed whose limit has rank-3. In contrast, the best low-rank approximation problem is well-posed for nonnegative tensors qi2016 . Given these phenomena for general tensors, there is a technical challenge in determining whether a new gauge norm as defined above can act as a convex surrogate for the rank of tensors.
@@ 2 Preliminaries
By definition, a rank-1 tensor can be written as the tensor product of vectors, that is $\psi=\bigotimes_{k=1}^{p}\theta^{(k)}$, where $\theta^{(k)}\in\mathbb{R}^{r_{k}}$. Equivalently, each tensor entry is $\psi_{x}=\prod_{k=1}^{p}\theta^{(k)}_{x_{k}}$, where $\theta^{(k)}_{x_{k}}$ is the $x_{k}$-th element of vector $\theta^{(k)}$ for any index value $x_{k}\in[r_{k}]$. When obvious, we will use $\theta_{x_{k}}$ instead of $\theta^{(k)}_{x_{k}}$.
@@
$rank(\psi)=\min\{q\ |\ \psi=\sum_{k=1}^{q}\psi^{k},\psi^{k}\in\mathcal{B}_{\infty}\text{ for }k\in[q]\}.$
Using this definition of tensor rank, a cp decomposition of the tensor is given by $\psi=\textstyle\sum_{k=1}^{rank(\psi)}\psi^{k}$.
@@
where the constraint uses a new norm $\|\psi\|_{\pm}$ that we design below. We show that our norm does in fact act as a convex surrogate for tensor rank, analyze its complexity, and demonstrate that it leads to a practical global algorithm for the tensor completion problem.
@@ 3 Gauge Norm for Tensors
We construct a norm for general tensors using a gauge function chandrasekaran2012convex ; Jaggi_2013 ; bugg2022nonnegative . To start with, consider rank-1 tensors. Let $\mathcal{S}_{\lambda}$ be the set of rank-1 tensors such that each entry of the tensor has an absolute value of some $\lambda\in\mathbb{R}_{+}$:
@@
Note that the difference between the sets $\mathcal{S}_{\lambda}$ and $\mathcal{B}_{\lambda}$ is that their rank-1 tensors have entries in $\{-\lambda,\lambda\}$ and $[-\lambda,\lambda]$ respectively. Our first step in defining a norm is to relate the convex hull of these sets of rank-1 tensors.
Proposition 1
The convex hulls of these sets are the same, meaning we have $\mathcal{C}_{\lambda}:=conv(\mathcal{B}_{\lambda})=conv(\mathcal{S}_{\lambda})$.
Proof
The proof for this proposition is similar to the proof of Proposition 2.1 of bugg2022nonnegative , but where a multilinear optimization problem is formulated by restricting the entries to $[-1,1]$ instead of $[0,1]$ as in (4).
[…]
$\max$ $\langle\varphi,\psi\rangle$ (4)
[…]
Proposition 2.1 of drenick1992multilinear shows there exists a global optimum $\psi^{\prime\prime}$ of (4) with $\psi^{\prime\prime}\in\mathcal{S}_{\lambda}$. By construction, we must have $\langle\varphi,\psi^{\prime\prime}\rangle\geq\langle\varphi,\psi^{\prime}\rangle$, which implies $\langle\varphi,\psi^{\prime\prime}\rangle\geq\langle\varphi,\psi\rangle+\delta$ for all $\psi\in conv(\mathcal{S}_{\lambda})$. But this last statement is a contradiction since $\psi^{\prime\prime}\in\mathcal{S}_{\lambda}\subseteq conv(\mathcal{S}_{\lambda})$.
Remark 1
@@ 3.1 Defining the Norm
Proof
We use the result from Example 3.50 of rockafellar2009 to conclude that $\|\cdot\|_{\pm}$ is a norm. To apply the result, we check that the required conditions on $\mathcal{C}_{1}$ hold.
@@
The final, non-trivial condition required is that $\mathcal{C}_{1}$ has a non-empty interior. To prove this, we use Theorem 2.4 of Rockafellar_2015 that the dimension of $\mathcal{C}_{1}$ is the maximum of dimensions of simplices included in it. We construct a simplex in $\mathcal{C}_{1}$ that has dimension $\pi=\Pi_{i}r_{i}$, and hence $\mathcal{C}_{1}$ has a dimension of at least $\pi$. But the flattened vector space of tensors also has dimension $\pi$; consequently, the dimension of $\mathcal{C}_{1}$ is at most $\pi$. Hence, $\mathcal{C}_{1}$ has a dimension $\pi$, which is the full dimension of the space, implying that the set $\mathcal{C}_{1}$ must have a non-empty interior.
The rest of the proof constructs such a simplex of dimension $\pi$. Consider a polytope $D=conv(0\>\cup\>\{d^{x}\}_{x\in\mathcal{R}})$. Here, for any $x=(x_{1},\cdots,x_{p})\in\mathcal{R}$, consider $d^{x}=\bigotimes_{k=1}^{p}\beta^{x_{k}}$ with each vector $\beta^{x_{k}}\in\mathbb{R}^{r_{k}}$ defined as
@@
With this, $\mathcal{C}_{1}$ is shown to satisfy all the required conditions, which means that the proposition holds.
Towards our end goal of developing a practical algorithm for tensor completion using this gauge norm, our next result provides an alternative characterization of the vertices of $\mathcal{C}_{\lambda}$. This result is important because it shows that these vertices can be represented by linear inequality constraints using (binary) integer variables. This is unlike, say, the nuclear norm, which under the $\ell_{2}$ norm defines an inherently nonlinear (continuous) feasible region.
Proposition 3
@@
Proof
We show that the constraints defining the sets $\mathcal{S}_{\lambda}$ and $\widehat{\mathcal{S}}_{\lambda}$ are equivalent. Consider some $x\in\mathcal{R}$. From the definition of $\mathcal{S}_{\lambda}$ in (3), we have $\psi_{x}=\lambda\prod_{i=1}^{p}\theta_{x_{k}}$. Define $y_{x,k}=\prod_{i=k}^{p}\theta_{x_{k}}$, for $k\in[p]$ so that $\psi_{x}=\lambda y_{x,1}$. Or equivalently, in a recursive relationship, $y_{x,k}=\theta_{x_{k}}y_{x,k+1}$ for $k\in[p-1]$ and $y_{x,p}=\theta_{x_{p}}$. The recursive constraints can be thought of as a negated-XOR relation, and linearized by transformations for conjunctive and disjunctive statements (see Section 2.5 of Conforti_Cornuéjols_Zambelli_2014 ). These linearized constraints correspond to constraints 2-5 in the definition of $\widehat{\mathcal{S}}_{\lambda}$ as given above. This shows that for each $x\in\mathcal{R}$, the tensor $\psi_{x}$ is defined the same in both $\mathcal{S}_{\lambda}$ and $\widehat{\mathcal{S}}_{\lambda}$.